OpenStudy (little.cherry23):

help please !! 1.)How many solutions does the system of equations have 15y-3x+9 y=-1/5x+9/15 a.)one** b.)two c.)none d.)infinitely many 2.) what is the slope of the line that passes through the pairs of points ? (-2,1),(2,17) a.)4/17 b.)-15 c.)-4** d.)4 @Qwertty123

OpenStudy (qwertty123):

Yes hello!

OpenStudy (little.cherry23):

helloo

OpenStudy (qwertty123):

Guess what another link just for you! It is really easy once you got it! https://www.desmos.com/calculator there you go!

OpenStudy (little.cherry23):

ive tried that link and it dont help as much

OpenStudy (qwertty123):

did you put in the two equations in two different lines?

OpenStudy (qwertty123):

@Little.cherry23

OpenStudy (little.cherry23):

yes id did i just dont clearly under stand how to use that

OpenStudy (qwertty123):

oh well you put in the equations then you hover where the two lines intersect... okay?

OpenStudy (anonymous):

i think the answers are b and d

OpenStudy (little.cherry23):

and how did you get that

OpenStudy (qwertty123):

@Improvement Don't give direct answers!

OpenStudy (danjs):

a quick check you can always do, -any multiple of a linear equation is the same line, like this x + 2y = 1 5x + 10y = 5 same lines

OpenStudy (little.cherry23):

so its b and d ?

OpenStudy (little.cherry23):

ive been at this for 6 hours

OpenStudy (qwertty123):

He is so correct! @DanJS

OpenStudy (danjs):

there is a quick orthogonal test with the dot product

OpenStudy (little.cherry23):

?

OpenStudy (qwertty123):

What?

OpenStudy (little.cherry23):

also how would you graph >5x-2 or <5y-2 ??

OpenStudy (danjs):

15y = 3x+9 y=-1/5x+9/15 i just mentioned that check, because here, you can notice the second equation is just 1/15 of the first, a multiple of each other, same lines, overlap infinitly many times

OpenStudy (danjs):

graph it the same as an equality, except dotted lines for > and < if not also equal to, then shade the right area according the the inequality

OpenStudy (anonymous):

i am correct

OpenStudy (anonymous):

can i have a medal plz i need some

OpenStudy (danjs):

two lines can not intersect twice, there will always be -1 unique solution at the intersection (x,y) - 0 solutions, for parallel non intersecting lines - infinity , for a system like this prob where they are the same overlapping line

OpenStudy (qwertty123):

Your correct for what @Improvement

OpenStudy (anonymous):

i knew i was

OpenStudy (danjs):

solving the system for parallel or same lines will not result in any variable values for x or y, usually, if you get a false statemtnt, like 5 = 8, then it is no solution, parallel if you get a true statemtn like 5=5, then same lines, infinity solutions

OpenStudy (qwertty123):

I knew I was.... what? @Improvement

OpenStudy (danjs):

b and d, , the first is not 'two'

OpenStudy (anonymous):

i think it is

OpenStudy (anonymous):

dan just think

OpenStudy (qwertty123):

I think she got the answers already! :)

OpenStudy (danjs):

a system of two linear equations will not have two solutions, 1,0,or infinity

OpenStudy (danjs):

draw two lines and have them intersecting twice, then i will believe it

OpenStudy (qwertty123):

What number are you on @DanJS

OpenStudy (little.cherry23):

how do you put -4 >32 on a line graph @im

OpenStudy (anonymous):

wait dan let me check again

OpenStudy (danjs):

it looks like everyone is saying two for the first prob, they are the same lines, infinity is the correct choice

OpenStudy (anonymous):

lets see mabey im wrong

OpenStudy (anonymous):

wait a sec

OpenStudy (qwertty123):

You cant @Little.cherry23 that line can't be graphed! :/

OpenStudy (danjs):

any x and y value will work in both of them

OpenStudy (little.cherry23):

x=-4/5y+3 4x-5y=-15 a.)parallel b.)perpendicular c.)neither

OpenStudy (qwertty123):

That one you can do! You can graph then both!

OpenStudy (little.cherry23):

im not sure if the x and y values are right because alot of them are cut out my questions so its really hard to answer my questions

OpenStudy (anonymous):

yeah i think dan is right this time

OpenStudy (qwertty123):

Hey Improvement how many medals do you have?

OpenStudy (anonymous):

not many ussually dont help i just started

OpenStudy (anonymous):

help me get some???

OpenStudy (little.cherry23):

dan is the very top two questions d and d ?

OpenStudy (qwertty123):

I will give you a medal @Improvement for trying to help her out! :D

OpenStudy (anonymous):

yes its d and d

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

can everyone in here give me a medal lol

OpenStudy (qwertty123):

no @Improvement ! :) I medal you can only give away. I think @Little.cherry23 gave hers away.

OpenStudy (danjs):

yes, for this other one x = -4/5y + 3 4x - 5y = -15 -------------- you can compare slopes in the y=m*x+b form, same is parallel negative recipricle is perpendicular, or both in standard form 5x + 4y = 15 4x - 5y = -15 ----------- notice 5*4 + (4*-5) = 0, if you do that pattern and get zero, they are perpendicular (vector dot product being zero) but forget that

OpenStudy (qwertty123):

So true!

OpenStudy (danjs):

<a,b>*<c*d>, direction vectors , dot product is zero if they are perpendicular from standard form for the lines Ax + By =C just multiply the two x coefficients and add that to the product of the two y coefficients, quick check to say perpendicular lines if you get a zero,

OpenStudy (anonymous):

dan can u answer a couple of my questions